OpenStudy (anonymous):

which of the followingis a solutuion to the inequality y>3x-4? a)(2,2) b)(-1,-9) c)(1,1) d)(0,-6)

OpenStudy (anonymous):

Did you know that if you plug in the numbers, you could just check. (x,y) = (.. , .. ). Throw them into the inequality and see which ones work. If you end up with something like 3 > 4, then that one didn't work...

OpenStudy (anonymous):

yea i dont even no how to do the problem....

OpenStudy (anonymous):

Take (b) for example. it's given you (-1, -9). This is saying that it wants you to try x = -1, and y = -9. Now have a look at the inequality, with those numbers for x and y. y = -9. 3x-4 = 4 * (-1) - 4 = -4 - 4 = -8. Now, is y > 3x-4? No. So (b) is not a possible solution. Do the same with the rest.

OpenStudy (anonymous):

can i just have the answer bec i have 30 mins left t take this timed test an i have a whole bunch more problems to do..

OpenStudy (anonymous):

In the time it took you to type that, you could have solved it yourself, rather than thinking that a stranger might do it all for you for some unknown reason.

OpenStudy (anonymous):

Honestly i dont get how your teaching it. I was learned another way and i forgot it as well.